OpenStudy (amtran_bus):

.4 percent of 4 percent of 1.25

OpenStudy (amtran_bus):

I said .004*1.25 = .005 or 1.245 Then 1.245 (.04) =.0498 or 1.1952 but that is wrong. Apparently you can say .004*.04*1.25 and get .0002. What concept am I missing or simple mathematical flaw am I making?

OpenStudy (kydawg):

How did you get 1.245 from 0.005?

OpenStudy (amtran_bus):

Don't you just subtract? Or am I just not seeing what you are saying?

OpenStudy (amtran_bus):

Nope. I see what i did.

OpenStudy (welshfella):

4% as a decimal is 0.04 so we have 0.04 * 0.04 * 1.25

OpenStudy (amtran_bus):

.005*.04 gives the answer

OpenStudy (amtran_bus):

So I can do it both ways, but it helps if I do my way correct lol

OpenStudy (kydawg):

You start the problem by finding 4% of 1.25. 0.04 * 1.25 = 0.05 Then you take 0.4% of that, so 0.004 * 0.05 = 0.0002.

OpenStudy (amtran_bus):

Yep. Thanks!

OpenStudy (welshfella):

oh sorry I misread the first part is 0.4 percent which is 0.004